A signal has a period T of 0.014 µs, state it's frequency (showing formula)
Answer #1

A signal has a period T of 0.014 µs; T=0.014 * 10-6 sec

The formula for frequency is: f (frequency) = 1 / T (period)

So, f (frequency) = 1/(0.014 * 10-6) = 71.42 * 106 Hz

Hence A signal has a period T of 0.014 µs, frequency would be (71.42 * 106 )Hz.

/*If this helps you, please let me know by giving a positive thumbs up. In case you have any queries, do let me know. I will revert back to you. Thank you!!*/